# A Closer Look at TabPFN v2: Strength, Limitation, and Extension

### v1

@@ Appendix B Evaluation Details
Appendix C Detailed Results
@@
  • We present the main results of TabPFN v2 on 300 datasets in LABEL:tab:main_results. The table includes accuracy for classification tasks and RMSE (Root Mean Squared Error) for regression tasks, along with the corresponding mean and standard deviation for each dataset. Notably, we excluded 27 datasets from these results in LABEL:tab:main_results, as they were used by TabPFN v2 to select the best checkpoint. These excluded datasets, which are not shown in Figure 2 and Figure 3, include:
@@
  • We demonstrate the performance of various models on 12 multi-class classification tasks with more than 10 classes in Table 7. The table provides the mean accuracy of models like KNN, TabPFN-v2*, XGBoost (XGB), CatBoost (CatB), Random Forest (RForest), ModernNCA (MNCA), MLP, Logistic Regression (LogReg), and RealMLP, showcasing how they perform on multi-class tasks with a larger number of classes. Additionally, we compare PFN-v2-ECOC, a multi-class classification solution provided by Hollmann et al. (2025). This method extends TabPFN-v2 by leveraging Error-Correcting Output Codes (ECOC) to enhance multi-class classification performance [^2] https://github.com/PriorLabs/tabpfn-community/blob/main/src/tabpfn_extensions/many_class/many_class_classifier.py.
@@
  • We present a comparison of FT-T (Gorishniy et al., 2021) and MNCA (Ye et al., 2025) with and without randomized feature tokens on tiny benchmark datasets from Ye et al. (2024b) in Table 9. The first two columns represent FT-T and MNCA with randomized feature tokens, while the last two show the original FT-T and MNCA methods. The table demonstrates the effect of randomized feature tokens on model performance, with accuracy for classification tasks and RMSE for regression tasks. FT-T and MNCA implementations, including their hyperparameter search space, follow the approach described in Liu et al. (2024), and the results were obtained after 100 rounds of hyperparameter tuning.
@@
CLL_SUB_111 62.90 72.46 57.39 70.43 73.04 70.14 73.91 70.14 57.68 73.17 65.51 71.59
BASEHOCK 96.31 97.01 71.88 97.46 95.29 96.73 96.99 69.09 97.41 97.47 96.32 95.87
Prostate_GE 81.27 86.98 80.00 87.94 89.52 87.94 91.43 95.24 88.57 94.60 84.13 94.92
PCMAC 88.21 88.53 66.48 90.15 91.64 92.20 87.15 92.70 90.76 86.86 88.21 92.01
@@
RELATHE 88.18 90.54 75.03 90.23 87.11 87.30 90.49 86.36 87.65 86.81 89.92 90.35
SMK_CAN_187 63.51 66.84 69.47 69.82 66.49 70.70 72.11 71.05 71.75 73.28 70.71 71.40
warpPIE10P 98.41 99.05 92.38 100.0 94.92 98.57 100.0 100.0 100.0 99.84 97.79 98.89
leukemia 90.22 95.11 86.67 94.67 97.78 92.00 96.00 92.44 93.33 95.33 94.00 94.22
orlraws10P 97.67 98.33 92.00 99.00 84.33 99.00 99.00 92.00 99.33 98.33 92.67 99.00
GLIOMA 58.00 60.67 68.00 67.33 66.67 64.00 64.00 62.67 69.33 68.00 69.91 66.67
warpAR10P 83.08 85.64 53.08 97.44 81.28 87.18 97.69 90.77 95.38 92.05 90.04 87.44
TOX_171 76.00 88.19 70.86 90.48 78.10 78.67 90.29 80.95 82.48 86.86 85.52 83.05
lung 91.54 95.45 93.66 95.28 93.66 92.68 95.12 95.28 93.50 95.45 95.43 93.01
ALLAML 87.56 95.56 81.33 96.89 96.00 96.44 92.00 92.89 93.78 92.44 91.14 94.67
colon 78.46 78.97 76.92 83.08 74.87 82.56 86.15 81.54 78.46 80.00 78.46 77.95
gisette 97.21 97.57 95.04 97.86 97.55 96.82 97.51 97.35 97.26 97.37 97.18 97.78
arcene 81.67 85.50 84.50 81.00 75.00 86.83 88.00 83.67 88.33 90.67 85.33 85.00
Mean 82.86 87.11 77.29 88.83 84.76 86.87 89.36 85.25 87.29 88.95 86.37 87.48
@@
Dataset KNN PFN-v2* PFN-v2-ECOC XGB CatB RForest MNCA MLP LogReg RealMLP
100-plants-texture 79.69 90.94 84.92 77.06 89.73 82.65 80.52 83.92 86.88 88.35
100-plants-margin 77.50 88.56 79.40 74.25 84.06 82.79 77.60 80.44 79.69 83.58
100-plants-shape 60.31 79.52 63.38 56.15 65.19 64.33 70.10 47.33 65.94 72.08
UJI_Pen_Characters 36.26 45.71 44.20 30.35 38.88 34.24 44.03 37.75 19.41 46.37
texture 98.45 100.0 100.0 98.55 99.13 96.76 99.68 99.40 99.64 99.95
letter 94.90 97.57 97.78 96.26 96.75 91.56 97.96 96.40 75.80 98.31
walking-activity 60.29 61.22 61.92 65.06 64.92 61.74 64.85 60.64 27.02 65.13
helena 28.94 33.31 19.20 32.42 37.90 33.91 36.58 37.91 33.40 38.55
internet_usage 30.17 54.34 50.86 51.08 37.90 33.91 52.09 43.00 37.73 52.23
kropt 71.22 77.96 77.11 86.95 79.26 71.77 78.27 64.45 28.08 92.03
kr-vs-k 70.78 78.46 76.29 87.26 74.81 71.60 76.83 65.03 28.03 91.85
ASP-POTASSCO 34.75 43.50 45.27 42.24 41.08 42.86 37.45 29.63 35.14 41.70
Mean 61.94 70.93 66.69 66.47 67.47 64.01 68.00 62.16 51.40 72.51

### v2

@@ Appendix A Evaluation Details
Table 4: Explanation of the feature name abbreviations in the bank dataset used in Figure 4.
Feature Description
age Age of the client.
bal Average yearly balance in euros.
day Last contact day of the month.
dur Last contact duration in seconds.
camp Number of contacts performed during this campaign.
pdays Number of days since the client was last contacted (999 means never contacted).
prev Number of contacts performed before this campaign.
job Job type of the client (e.g., admin., technician, unemployed).
mar Marital status of the client.
edu Education level of the client.
def Whether the client has credit in default.
house Whether the client has a housing loan.
loan Whether the client has a personal loan.
cont Communication type used for the last contact (e.g., cellular, telephone).
mon Last contact month of the year.
pout Outcome of the previous marketing campaign.
Table 5: Explanation of the feature name abbreviations in the second dataset used in Figure 4.
Feature Description
state The U.S. state associated with the customer’s account.
al The number of days or months the account has been active.
pn The customer’s phone number (may serve as an identifier).
nvmm The number of voicemail messages received.
tdm Total number of minutes the customer talked during the day.
tdc Total number of calls made during the day.
tdc2 Total charges incurred during the day.
tem Total number of minutes talked during the evening.
tec Total number of evening calls made.
tec2 Total charges incurred during the evening.
tnm Total number of minutes talked at night.
tnc Total number of night calls made.
tnc2 Total charges incurred at night.
tim Total number of minutes talked internationally.
tic Total number of international calls made.
tic2 Total charges for international calls.
ac The area code of the customer’s phone number.
ip Whether the customer has an international calling plan (yes/no).
vmp Whether the customer has a voicemail plan (yes/no).
ncs Number of times the customer contacted customer service.
Appendix B Details of Comparison Baselines
This section provides implementation details of the TabPFN v2 variants used in our analysis for high-dimensional, many-class, and large-scale tasks (cf. Figure 6).
High-Dimensional Classification. To address high-dimensional tasks, we introduce a baseline variant, TabPFN v2-PCA, which incorporates dimensionality reduction. Specifically, TabPFN v2-PCA reduces the feature dimension to 500 using PCA to satisfy the input constraints of TabPFN v2. This process is repeated multiple times with different PCA projections, and the resulting predictions are aggregated via bagging to improve robustness.
Many-Class Classification. We consider the following variants of TabPFN v2 to address the 10-class limit in classification tasks:
  • TabPFN v2-ECOC: We use the implementation provided in the official TabPFN extensions repository, which applies Error-Correcting Output Codes (ECOC).
  • TabPFN v2-DPT: We encode each class label as a $t$-digit decimal string and train a separate TabPFN v2 to predict each digit. For instance, a 15-class problem is decomposed into two subproblems: one for the tens digit (classes $\{0,1\}$) and one for the ones digit (classes $\{0,\dots,9\}$). The predicted digits are then decoded to recover the final class label. The same strategy is also proposed in [49].
We implement TabPFN v2 based on TabPFN v2-DPT for efficiency. Specifically, TabPFN v2 permutes the class-to-digit mapping $\sqrt{C}$ times. For fair comparison, we also increase the number of ensembles in TabPFN v2-DPT to $\sqrt{C}$ per digit to match the total number of predictions.
Large-Scale Tasks. As the size of training data increases, the benefit of pre-trained knowledge may diminish. We consider the following variants of TabPFN v2 to scale to larger datasets:
  • TabPFN v2-DT: A shallow decision tree is trained with a minimum split size of 10,000. The tree partitions the dataset into smaller subsets, and a separate TabPFN v2 model is applied to each leaf node. At inference, a test instance is routed through the tree to a corresponding leaf, where it is predicted by the respective TabPFN v2 model.
  • TabPFN v2-B: A bagging-based variant that randomly samples 10,000 training examples four times and aggregates their predictions.
  • TabPFN v2-K: Selects a representative subset of 10,000 training examples based on proximity to KMeans-derived prototypes.
Our TabPFN v2 -DF is an extension of TabPFN v2-DT to a forest-based ensemble. Specifically, we sample 32 subsets from the original training data, each containing 60% of the samples (without replacement), and train a separate TabPFN v2-DT model on each subset. During inference, predictions from all 32 models are aggregated—e.g., by majority voting or averaging—depending on the task type.
Appendix C Additional Feature Extraction Results
In this section, we provide further results to validate the effectiveness of our leave-one-fold-out strategy for feature extraction with TabPFN v2. We also compare this supervised approach to two unsupervised embedding extraction methods.
C.1 Additional Results for Figure 5
Figure 7 visualizes instance embeddings extracted using our leave-one-fold-out strategy on two datasets: website_phishing (first row, three classes) and KDD (second row, binary). Training and test examples are shown as darker crosses and lighter circles, respectively, and colored by class. These visualizations support our claim that TabPFN v2, when combined with our proposed strategy, can serve as an effective feature encoder.
(a) Raw Feature
(b) Vanilla
(c) Layer 3
(d) Layer 6
(e) Layer 9
(f) Layer 12
Figure 7: Visualization of the extracted instance features from two datasets: website_phishing (first row, three classes) and KDD (second row, binary). Colors indicate classes, where Darker crosses and lighter circles denote training and test examples, respectively. (a) shows the raw input features (e.g., ${\bm{x}}_{i}$), while (b) presents embeddings from the vanilla strategy. (c)-(f) display embeddings produced by our method at different layers. Classification accuracy is reported by training a linear logistic regression model on the training embeddings and evaluating on the test set.
Table 6: Average rank (lower is better) of TabPFN v2 and a linear classifier trained on the extracted embeddings across 29 classification datasets. In addition to the supervised feature extraction strategy considered in Section 6 (including the vanilla one, our leave-one-fold-out, and the version based on the combined features), we compare with two unsupervised embedding extraction approaches by appending a column of dummy labels with zero values and permuting each column as labels, respectively.
$\downarrow$ TabPFN v2 Vanilla Dummy Permute Ours Combined
Rank 2.66 5.72 4.90 3.69 2.16 1.88
C.2 Comparison Between Supervised and Unsupervised Feature Extraction
Our leave-one-fold-out strategy explicitly incorporates label information and accounts for the distinct roles of training and test instances. To further understand the nature of the extracted embeddings, we compare this supervised strategy to two unsupervised alternatives provided by the TabPFN extensions repository.
Let $X\in\mathbb{R}^{n\times d}$ denote a dataset with $n$ samples and $d$ features:
  • Unsupervised-Dummy: Each sample is paired with a constant pseudo-target $\mathbf{y}=\mathbf{0}\in\mathbb{R}^{n}$, forming a regression task. The embedding $E_{\text{D}}\in\mathbb{R}^{n\times h}$ is obtained by extracting the output tokens of the TabPFN regressor.
  • Unsupervised-Permute: For each feature $j\in\{1,\ldots,d\}$, we treat $\mathbf{x}^{(j)}=X_{:,j}$ as a pseudo-target and use the remaining features $X^{(-j)}$ as input. Depending on the type of $\mathbf{x}^{(j)}$, TabPFN is applied in classification or regression mode to obtain $E^{(j)}\in\mathbb{R}^{n\times h}$. These embeddings are concatenated to a high-dimensional form:
    $E_{\text{P}}=\text{concat}(E^{(1)},E^{(2)},\ldots,E^{(d)})\in\mathbb{R}^{n\times(d\cdot h)}.$
We compare these unsupervised methods with our supervised leave-one-fold-out strategy in Table 6. Overall, unsupervised approaches underperform compared to the supervised ones and fail to recover the classification ability of TabPFN v2. This performance gap arises because label information is introduced only post hoc via a linear classifier rather than during embedding extraction. Among the unsupervised methods, the permutation-based approach performs better, likely due to its ability to encode attribute-specific structure.
Figure 8 presents a visual comparison of embeddings produced by the three methods using the same color and marker scheme as in Figure 7. Since unsupervised methods lack label supervision during embedding generation, their embeddings tend to scatter broadly without forming well-separated clusters by class. These results further highlight the contrasting goals of supervised and unsupervised strategies—class separation versus feature distribution coverage, respectively.
(a) Raw Feature
(b) Vanilla
(c) Dummy
(d) Permute
(e) Ours
Figure 8: Comparison between unsupervised and supervised (ours) feature extraction. Visualization of extracted embeddings for four datasets: churn (first row, two classes), bank (second row, two classes), KDD (third row, two classes), and website_phishing (fourth row, three classes). We use crosses to denote training examples and circles to denote test examples. (a) shows the raw features, while (b) presents the embeddings extracted using the vanilla strategy. (c) and (d) refer to unsupervised embedding extraction approaches by appending a column of dummy labels with zero values and permuting each column as labels, respectively. (e) depicts the embeddings obtained using our proposed methods. The accuracy value is calculated by training a linear model (logistic regression) over the extracted embeddings on the training set and predicting on the test set.
Appendix D Influence of Key Modules and Meta-Feature Analysis
We investigate the influence of two key components in TabPFN v2, i.e., the feature engineering that pre-processes the raw features of a given tabular dataset and the post-hoc ensembleing. In addition, we analyze the conditions under which TabPFN v2 performs well or poorly through a meta-feature-based classification analysis.
Feature engineering. TabPFN v2 pre-processes the features of a given tabular dataset with various strategies, such as quantile, category shuffling, SVD, and power transform. Specifically, we examine the effects of adding fingerprint features (add_fingerprint_feature) and polynomial features (polynomial_features) to the raw tabular data. The results indicate that TabPFN v2 performs well even without the use of these engineered features, suggesting that, for the benchmark datasets of [76], these specific feature engineering techniques do not provide a significant improvement. This finding highlights the robustness of TabPFN v2 and its ability to handle raw features effectively, without the need for extensive pre-processing or feature construction. We show the influence of this step in Figure 9.
Model ensemble. Post hoc ensembling (PHE) involves applying TabPFN v2 to the datasets multiple times with different perturbations and aggregating the predictions of these base models at different temperatures. We show the change of performance of TabPFN v2 w.r.t. the number of ensemble numbers (i.e., the number of base models) in Figure 10. On the benchmark of [76], we observe that, overall, ensemble methods improve performance, with larger ensemble sizes yielding better results. However, we also note that even without ensembling, TabPFN v2 performs exceptionally well, and the relative performance gain from ensembling is limited. This suggests that while ensembling can provide further improvements, the base TabPFN v2 model is already highly effective on its own. The equivariant property described in [3] provides insight into this phenomenon. Since TabPFN v2 introduces random tokens to handle heterogeneous features, the model becomes less sensitive to the arbitrary ordering of features, effectively enforcing equivariance in this aspect. As a result, the benefits of ensembling through feature order permutations are less pronounced compared to TabPFN v1.
Meta-Feature Analysis of TabPFN v2 Performance. To better understand the conditions under which TabPFN v2 performs well or poorly, we conducted a meta-learning-based classification analysis using 300 datasets. Specifically, we used the average rank of TabPFN v2 across datasets as a performance indicator. The threshold for classification was set at the mean rank, 6.31. Datasets where TabPFN v2 achieved a rank lower than or equal to 6.31 were labeled as “Good”, while those with a higher rank were labeled as “Bad”. We extracted meta-features from each dataset and used them to train a decision tree classifier, aiming to distinguish between the “Good” and “Bad” categories. All the meta-features utilized in this task are detailed in Table 7, accompanied by their respective explanations. A visual depiction of a simple depth-3 decision tree is shown in Figure 11, and the tree reveals key factors that influence the effectiveness of TabPFN v2. The decision tree visualizes how to predict whether TabPFN v2 performs well (“Good”) or poorly (“Bad”) on a given dataset, based on dataset meta-features: The root node splits on the number of instances (nr_inst), indicating that TabPFN v2 tends to perform better on datasets with fewer than 24,350 samples. For these smaller datasets, the left subtree further splits on the mean joint entropy (joint_ent.mean), where higher values (greater than 3.028) are associated with improved performance. For datasets with lower mean joint entropy ($\leq$ 3.028), TabPFN v2 also tends to perform well when the number of rows is relatively small ($\leq$ 4,862). In contrast, the right subtree, which represents larger datasets, reveals that a low standard deviation of the interquartile range (iq_range.std) across features ($\leq$ 0.657) is linked to poorer model performance.
Figure 11: Decision tree for predicting TabPFN v2 performance (Good vs. Bad) based on dataset characteristics, constructed from experiments on 300 datasets. The tree splits on meta-features such as number of instances (nr_inst), joint entropy (joint_ent.mean), number of numerical features (nr_num), distance between minority and majority classes’ center of mass (gravity) and interquartile range (iq_range) statistics. Each split is chosen to maximize information gain. The leaf nodes indicate the predicted performance class, the Gini impurity, and class distribution. This tree provides insights into the types of datasets where TabPFN v2 is expected to perform well.
Table 7: Meta-features used in the meta-feature analysis of TabPFN v2 performance.
Meta-Feature Explanation
attr_conc The concentration coef. of each pair of distinct attributes.
class_conc The concentration coefficient between each attribute and class.
class_ent The target attribute Shannon’s entropy.
inst_to_attr The ratio between the number of instances and attributes.
mean The mean value of each attribute.
sd The standard deviation of each attribute.
var The variance of each attribute.
range The range (max - min) of each attribute.
iq_range The interquartile range (IQR) of each attribute.
nr_attr The total number of attributes.
sparsity The (possibly normalized) sparsity metric for each attribute.
t_mean The trimmed mean of each attribute.
nr_bin The number of binary attributes.
nr_cat The number of categorical attributes.
nr_num The number of numeric features.
nr_norm The number of attributes normally distributed based in a given method.
nr_cor_attr The number of distinct highly correlated pair of attributes.
gravity The distance between minority and majority classes’ center of mass.
nr_class The number of distinct classes.
joint_ent The joint entropy between each attribute and class.
attr_ent Shannon’s entropy for each predictive attribute.
cov The absolute value of the covariance of distinct dataset attribute pairs.
eigenvalues The eigenvalues of covariance matrix from dataset.
eq_num_attr The number of attributes equivalent for a predictive task.
max The maximum value from each attribute.
min The minimum value from each attribute.
median The median value from each attribute.
freq_class The relative frequency of each distinct class.
mad The Median Absolute Deviation (MAD) adjusted by a factor.
mad The Median Absolute Deviation (MAD) adjusted by a factor.
mut_inf The mutual information between each attribute and target.
nr_inst The number of instances (rows) in the dataset.
nr_outliers The number of attributes with at least one outlier value.
ns_ratio The noisiness of attributes.
imblance_ratio The ratio of the number of instances in the minority to the majority class.
attr_to_inst The ratio between the number of attributes.
Appendix E Detailed Results
@@
  • We present the main results of TabPFN v2 on 300 datasets in LABEL:tab:main_results. The table includes accuracy for classification tasks and RMSE (Root Mean Squared Error) for regression tasks, along with the corresponding mean and standard deviation for each dataset. Notably, we excluded 27 datasets from these results in LABEL:tab:main_results, as they were used by TabPFN v2 to select the best checkpoint. These excluded datasets, which are not shown in Figure 1 (right) and Figure 3, include:
@@
  • We demonstrate the performance of various models on 12 multi-class classification tasks with more than 10 classes in Table 10. The table provides the mean accuracy of models like KNN, TabPFN-v2*, XGBoost (XGB), CatBoost (CatB), Random Forest (RForest), ModernNCA (MNCA), MLP, Logistic Regression (LogReg), and RealMLP, showcasing how they perform on multi-class tasks with a larger number of classes. Additionally, we compare PFN-v2-ECOC, a multi-class classification solution provided by [29]. This method extends TabPFN-v2 by leveraging Error-Correcting Output Codes (ECOC) to enhance multi-class classification performance. [^3] https://github.com/PriorLabs/tabpfn-community/blob/main/src/tabpfn_extensions/many_class/many_class_classifier.py
@@
CLL_SUB_111 62.90 72.46 57.39 70.43 73.04 70.14 73.91 70.14 57.68 71.59 65.51 71.59
BASEHOCK 96.31 97.01 71.88 97.46 95.29 96.73 96.99 69.09 97.41 97.36 96.32 95.87
Prostate_GE 81.27 86.98 80.00 87.94 89.52 87.94 91.43 95.24 88.57 94.29 84.13 94.92
PCMAC 88.21 88.53 66.48 90.15 91.64 92.20 87.15 92.70 90.76 90.14 88.21 92.01
@@
RELATHE 88.18 90.54 75.03 90.23 87.11 87.30 90.49 86.36 87.65 89.95 89.92 90.35
SMK_CAN_187 63.51 66.84 69.47 69.82 66.49 70.70 72.11 71.05 71.75 72.10 70.71 71.40
warpPIE10P 98.41 99.05 92.38 100.0 94.92 98.57 100.0 100.0 100.0 100.0 97.79 98.89
leukemia 90.22 95.11 86.67 94.67 97.78 92.00 96.00 92.44 93.33 96.00 94.00 94.22
orlraws10P 97.67 98.33 92.00 99.00 84.33 99.00 99.00 92.00 99.33 99.67 92.67 99.00
GLIOMA 58.00 60.67 68.00 67.33 66.67 64.00 64.00 62.67 69.33 68.67 69.91 66.67
warpAR10P 83.08 85.64 53.08 97.44 81.28 87.18 97.69 90.77 95.38 96.67 90.04 87.44
TOX_171 76.00 88.19 70.86 90.48 78.10 78.67 90.29 80.95 82.48 87.24 85.52 83.05
lung 91.54 95.45 93.66 95.28 93.66 92.68 95.12 95.28 93.50 95.61 95.43 93.01
ALLAML 87.56 95.56 81.33 96.89 96.00 96.44 92.00 92.89 93.78 94.67 91.14 94.67
colon 78.46 78.97 76.92 83.08 74.87 82.56 86.15 81.54 78.46 79.49 78.46 77.95
gisette 97.21 97.57 95.04 97.86 97.55 96.82 97.51 97.35 97.26 97.23 97.18 97.78
arcene 81.67 85.50 84.50 81.00 75.00 86.83 88.00 83.67 88.33 92.00 85.33 85.00
Mean 82.86 87.11 77.29 88.83 84.76 86.87 89.36 85.25 87.29 89.73 86.37 87.48
@@
Dataset KNN PFN-v2* PFN-v2-DPT PFN-v2-ECOC XGB CatB RForest MNCA MLP LogReg RealMLP
100-plants-texture 79.69 90.94 82.67 84.92 77.06 89.73 82.65 80.52 83.92 86.88 88.35
100-plants-margin 77.50 88.56 81.94 79.40 74.25 84.06 82.79 77.60 80.44 79.69 83.58
100-plants-shape 60.31 79.52 72.15 63.38 56.15 65.19 64.33 70.10 47.33 65.94 72.08
UJI_Pen_Characters 36.26 45.71 33.38 44.20 30.35 38.88 34.24 44.03 37.75 19.41 46.37
texture 98.45 100.0 99.98 100.0 98.55 99.13 96.76 99.68 99.40 99.64 99.95
letter 94.90 97.57 96.69 97.78 96.26 96.75 91.56 97.96 96.40 75.80 98.31
walking-activity 60.29 61.22 57.28 61.92 65.06 64.92 61.74 64.85 60.64 27.02 65.13
helena 28.94 33.31 28.54 19.20 32.42 37.90 33.91 36.58 37.91 33.40 38.55
internet_usage 30.17 54.34 50.51 50.86 51.08 37.90 33.91 52.09 43.00 37.73 52.23
kropt 71.22 77.96 71.44 77.11 86.95 79.26 71.77 78.27 64.45 28.08 92.03
kr-vs-k 70.78 78.46 71.54 76.29 87.26 74.81 71.60 76.83 65.03 28.03 91.85
ASP-POTASSCO 34.75 43.50 41.88 45.27 42.24 41.08 42.86 37.45 29.63 35.14 41.70
Mean 61.94 70.93 65.67 66.69 66.47 67.47 64.01 68.00 62.16 51.40 72.51